# Bayesian inference for the Markov-modulated Poisson process with an outcome process

### v1

@@ 6 Discussion
In this paper, we have considered a generalization of both the hidden Markov model and the Markov-modulated Poisson process (MMPP) to an MMPP with an outcome process, and we have proposed an exact Gibbs sampler to facilitate Bayesian inference. Our data model also incorporates an unobserved terminating event (death or moving away from the area).
@@
This advance is driven by the need to analyze longitudinal data and model an unobserved death so as to comprehend the progression of a disease processes. A critical decision revolves around whether observation times are considered non-informative or informative, on the state of the Markov chain, which relates to the disease severity. Neglecting to account for an informative observation process in the disease dynamics or to incorporate an unobserved terminating event, such as death, can potentially lead to biased parameter estimates. For the former, patients with more severe conditions or acute symptoms might visit their physicians more frequently than those with milder conditions or no symptoms. Furthermore, observed variables such as patients’ drug consumption or healthcare costs are influenced by the disease status, which remains unobserved in claims data. With the large amount of claim data and healthcare data available, there is a pressing need to work on appropriate statistical modelling framework for longitudinal healthcare studies. The proposed technique for inferring death prevents the bias that would arise if a model without death were employed and either the full time window or a shortened window only up to the final event time were used. Our approach offers a framework for analyzing complex healthcare data, providing insights into disease progression and patient health/death states that can inform clinical decision-making and healthcare interventions.
@@
Since our algorithm is a Gibbs sampler, it requires no tuning parameters and could be used “off the shelf” by practitioners. Simulation studies indicate the effectiveness of the Gibbs sampler and show how the event times contribute vital information, especially when the outcome process lacks a clear separation between the states. By integrating observation times and death into the model, we enhance our understanding of disease dynamics and healthcare utilization patterns, thus advancing the methodology for analyzing longitudinal data in healthcare research. These extensions enable a more robust and comprehensive analysis, offering insights that can inform clinical decision-making and improve patient care strategies.
@@
Although the proposed model possesses a flexible structure, there remains some scope for future research to explore manifold possibilities in analyzing irregularly spaced longitudinal data. While the current model categorizes patients into discrete states, which is appropriate for some chronic diseases thought to have discrete stages of progression, relaxing the assumption of discrete states could yield insights. Relaxing the assumption that patients evolve through discrete states of underlying disease progression, would allow estimation of a continuum of severity that may better represent the trajectory of COPD and other chronic diseases. Substantial advances in statistical methodology would be required to extend this approach to a continuous state space and to allow the latent state distribution to have an interpretation as a continuous index or score.

### v2

@@ 6 Discussion
In this paper, we have considered a generalization of both the hidden Markov model and the MMPP to an MMPP with an outcome process, and we have proposed an exact Gibbs sampler to facilitate Bayesian inference. Our data model also incorporates an unobserved terminating event (death or moving away from the area).
@@
This advance is driven by the need to analyze longitudinal data and model an unobserved death so as to comprehend the progression of a disease processes. A critical decision revolves around whether observation times are considered non-informative or informative, on the state of the Markov chain, which relates to the disease severity. Neglecting to account for an informative observation process in the disease dynamics or to incorporate an unobserved terminating event, such as death, can potentially lead to biased parameter estimates. With the large amount of claim data and healthcare data available, there is a pressing need to work on appropriate statistical modelling framework for longitudinal healthcare studies. The proposed technique for inferring death prevents the bias that would arise if a model without death were employed with a shortened observation window only up to the final event time. Our approach offers a framework for analyzing complex healthcare data, providing insights into disease progression and patient health/death states that can inform clinical decision-making and healthcare interventions.
@@
Since our algorithm is a Gibbs sampler, it requires no tuning parameters and could be used “off the shelf” by practitioners. Simulation studies indicate the effectiveness of the Gibbs sampler and show how the event times contribute vital information, especially when the outcome process lacks a clear separation between the states. By integrating observation times and death into the model, we enhance our understanding of disease dynamics and healthcare utilization patterns, thus advancing the methodology for analyzing longitudinal data in healthcare research. In our application, we imposed a constraint that the process can only progress forward since COPD gets progressively worse over time (Cleveland Clinic, 2024). However, one could make no restrictions on the allowed transitions in the state space. In the absence of other information, the hidden state would then be non-identifiable with all permutations of the state labels equivalent. To avoid the ensuing identifiability issue, meaningful priors could be placed on the outcome models for each state or the outcome process event rates – that is, we would distinguish the states via their impact on the observable quantities or imposing certain restrictions on $\mathbf{Q}$ (see Section 4 in Luo et al., 2021 for a detailed discussion).
@@
Although the proposed model possesses a flexible structure, there remains some scope for future research to explore several possibilities in analyzing such irregularly spaced longitudinal data. The current setup assumes time-homogeneous processes $X_{s}$ and $N_{s}$. Both processes may plausibly depend on time. For instance, within any given health state, older patients might engage more frequently with the healthcare system. However, introducing complex structures to model the transition in $X_{s}$ poses computational challenges in both frequentist and Bayesian inference (see for example, Kendall et al., 2024, and a discussion in the Supplementary Material of this article). In addition, the issue of identifying the number of hidden states in MMPPs has been a long-standing challenge. In the context of HMMs, this topic has been explored using approaches such as marginal likelihood, cross-validated likelihood, and complete-data likelihood (see, Celeux and Durand, 2008; Pohle et al., 2017; Luo and Stephens, 2021); however, model selection criteria based on these quantities become unreliable as the dimensionality increases beyond 50 (Chen and Chen, 2008). Moreover, while the current model categorizes patients into discrete states, which is appropriate for some chronic diseases thought to have discrete stages of progression, relaxing the assumption of discrete states could yield insights. Relaxing the assumption that patients evolve through discrete states of underlying disease progression, would allow estimation of a continuum of severity that may better represent the trajectory of COPD and other chronic diseases. Substantial advances in statistical methodology would be required to extend this approach to a continuous state space and to allow the latent state distribution to have an interpretation as a continuous index or score. For example, one could use the features included in comorbidity indices to measure multimorbidity in terms of the ability to predict future mortality and health services use. The inference problem for this type of assumption involves multivariate diffusion processes, using discrete-time data that may be incomplete and affected by measurement errors (e.g., Golightly and Sherlock, 2022).
@@ Acknowledgments
Supplementary Materials for “Bayesian inference for the Markov-modulated Poisson process with an outcome process”
Appendix A Using the Metropolis–Hastings algorithm for inference
We have presented an exact Gibbs sampler for inference on a Markov modulated Poisson process with an observation process at event times. If non-conjugate priors were used, one would need to use the Metropolis–Hastings algorithm. The Gibbs sampler has two appealing properties compared with the Metropolis–Hastings algorithm.
Firstly, it is automatic: the user can simply apply it to their data and the algorithm will run straight “out of the box”. By contrast, the Metropolis–Hastings algorithm requires, firstly a choice of blocking - should all the $\boldsymbol{\beta}$ parameters be updated simultaneously then all of the $\mathbf{Q}$ parameters and then the vector $\boldsymbol{\lambda}$, should all be updated at once or is some other choice preferable? Then, for the commonly used random walk proposal, for example, proposal scalings and variance matrices need to be chosen. These either require initial tuning runs or some form of adaptive MCMC (e.g. Roberts and Rosenthal, 2009) which in turn needs a choice of hyperparameters and careful watching.
Secondly, sampling from the conditional posterior explores the whole of the conditional state space for the parameter and with guaranteed acceptance. For example, with Gaussian observations, the covariate vector $\boldsymbol{\beta}$ is sampled directly from its conditional posterior whatever the dimension, $\mathsf{dim}(\boldsymbol{\beta})$. By contrast with a random walk Metropolis proposal the magnitude of the proposed jump in $\boldsymbol{\beta}$ must shrink in proportion to $1/\sqrt{\mathsf{dim}(\boldsymbol{\beta})}$ (e.g. Roberts and Rosenthal, 2001). The Metropolis-adjusted Langevin algorithm and Hamiltonian Monte Carlo have better scaling, but both require the computation of gradients of the log posterior.
Appendix B Non-homogeneous transitions
Inherent in the use of a time-homogeneous hidden Markov model is that the waiting times between the changes of state of the hidden process are exponentially distributed and that the probability of a particular change given that a state change has occurred does not depend on the amount of time already spent in the current state. However, it is possible that this is not the case in reality. We explain why inference using both the timing of the events, $\tau_{1},\tau_{2},\dots$ and the observations $o_{1},o_{2},\dots$ is difficult in the general scenario, but then illustrate a particular case where a natural extension of our algorithm is possible: when waiting times have an Erlang distribution.
It is helpful to re-examine why relatively tractable inference, as described in Sections 3 and 4, has been possible. Transitions of the underlying Markov model occur according to constant-rate Poisson processes. For example, in a chain with two live states and a death state. If $X_{0}=1$ then there are two Poisson processes, one with rate $q_{1,2}$ and one with rate $q_{1,3}$, and whichever has the first event determines the next change of state. Crucially, we have also modeled the events in the observation process as arising from a Poisson process; in our running example, since $X_{0}=1$, this has a rate of $\lambda_{1}$. The fact that both transitions and observation events occurred according to Poisson processes allowed us to consider an extended state space and write down the generator in Section 3:
$\begin{pmatrix}\mathbf{Q}-\boldsymbol{\Lambda}&\boldsymbol{\lambda}\\ \mathbf{0}&0\end{pmatrix}.$
Exponentiating the generator produced transition probabilities that enabled us to employ the forward-backward algorithm to simulate the states of the hidden chain at the event times of the observation process. Maximum-likelihood and Bayesian inference for homogeneous hidden Markov models is covered in depth in Cappé et al. (2005) and there have been many advances for the homogeneous case since. Methods for Bayesian inference on the Markov modulated Poisson process are described in Fearnhead and Sherlock (2006); Rao and Teh (2013); maximum-likelihood inference on a combined homogeneous HMM with an informative observation process is performed via the EM algorithm in Lange et al. (2015) and Mews et al. (2023). To the best of our knowledge this article is the first to address Bayesian inference the combination.
One natural mechanism for allowing for different waiting distributions and time-dependent transition probabilities is by making the Poisson processes for the transitions non-homogeneous. There has been work in the sphere of inference for non-parametric Poisson intensities for a general non-Markovian model where the states themselves are observed Titman (2011); Kendall et al. (2024); Putter and Spitoni (2018), maximum-likelihood for non-homogeneous hidden Markov models using numerical integration Titman (2023), and Bayesian estimation for a Poisson process with a non-parametric intensity Adams et al. (2009); Alie et al. (2023). Whilst the general technique of particle MCMC Andrieu et al. (2010) can always be employed as long as it is possible to simulate from the hidden process, it involves further choices such as the number of particles and the type of particle filter to be used. We would like to extend our relatively straightforward methodology to the non-homogeneous case, but there is no fixed rate matrix to exponentiate and so the transition probabilities between event times of the observation process are intractable (e.g., Kendall et al., 2024); moreover there would, in general, be no conjugate prior, so parameter inference would require the Metropolis–Hastings algorithm. We, therefore, see no natural extension of our method to the general case. However, it is possible to extend our method to the special case of a model where the timings for the potential state changes have Erlang distributions.
An Erlang distribution is a Gamma distribution with an integer shape parameter, $k$, and a rate parameter $q$. Hence, an $\mathsf{Erlang}(k,q)$ random variable is a sum of $k$ independent $\mathsf{Exp}(q)$ random variables. Transitions can, therefore be modeling by adding $k-1$ interim states for each actual state transition. We illustrate this by considering the three-state model from Section 5, where State 3 corresponds to death and where there are no transitions back to lower-numbered states. We allow for a single interim transition for each state change and label the states with the partial transitions that have occurred in brackets. We also allow the first partial transition towards death to occur from either State 1 or State 2, so that the marginal time to death does not have an Erlang distribution unless $q_{1,3}=q_{2,3}$; we note that in the Markov chain in Section 5, the marginal time to death does not have an exponential distribution unless $q_{1,3}=q_{2,3}$. The generator is
$\mathbf{Q}=\begin{array}[]{r|ccccccc|}&1&1(2)&1(3)&1(2,3)&2&2(3)&3\\ \hline\cr 1&-&q_{1,2}&q_{1,3}&0&0&0&0\\ 1(2)&0&-&0&q_{1,3}&q_{1,2}&0&0\\ 1(3)&0&0&-&q_{1,2}&0&0&q_{1,3}\\ 1(2,3)&0&0&0&-&0&q_{1,2}&q_{1,3}\\ 2&0&0&0&0&-&q_{1,3}&0\\ 2(3)&0&0&0&0&0&-&q_{2,3}\\ 3&0&0&0&0&0&0&0\end{array}~{},$
where for simplicity of presentation, the minus sign, --, in each row stands for the negative quantity that will make the entries in the row sum to 00. The corresponding vector for the rate of the observation process is $\boldsymbol{\lambda}=[\lambda_{1},\lambda_{1},\lambda_{1},\lambda_{1},\lambda_{2},\lambda_{2},0]^{\top}$.
Appendix C Continuous-path versus Discrete-time Likelihoods
In a CTHMM scenario it is possible to integrate out the hidden Markov chain between observation times to give the likelihood
$\mathcal{L}(\mathbf{Q};x_{[0,\tau]})=\prod_{t=1}^{T}\exp\left\{\mathbf{Q}\Delta_{t}\right\}_{x_{\tau_{t}},x_{\tau_{t+1}}},$
where $\Delta_{t}=\tau_{t+1}-\tau_{t}$ and $\tau_{T+1}=\tau_{e}$. That is, we impute the states at each observation time but we do not fill in the full continuous-time Markov chain between states. In our setting, this could be adapted to include the information that there were no observation events in each inter-event interval and that there were events at the observation times:
$\mathcal{L}(\mathbf{Q},\boldsymbol{\lambda};x_{[0,\tau]})=\left[\prod_{t=1}^{T-1}\exp\left\{(\mathbf{Q}-\boldsymbol{\Lambda})\Delta_{t}\right\}_{x_{\tau_{t}},x_{\tau_{t+1}}}\lambda_{x_{\tau_{t+1}}}\right]\exp\left\{(\mathbf{Q}-\boldsymbol{\Lambda})(\tau_{e}-\tau_{T})\right\}_{x_{\tau_{T}},x_{\tau_{e}}}$
as the transition probability. Only imputing the states at observation times would reduce the amount of computation a little and still permits a Gibbs step for updating the $\mathbf{B}$ matrix. As mentioned in the discussion, the evolution process occurs in continuous time and so it is natural to impute the full continuous-time process; however, there is also a more pragmatic reason to do this. If one only imputed the states at event times, one would be forced to perform inference using Metropolis—Hastings for both $\mathbf{Q}$ and $\boldsymbol{\lambda}$, and deal with the issues that we mention earlier in the supplementary material. The full underlying Markov chain is exactly what is needed for a Gibbs sampler on $\mathbf{Q}$ and $\boldsymbol{\lambda}$, and that is why our algorithm samples the full underlying Markov chain.
Appendix D Application: Additional results
The results from the four-state Hidden Markov model appear in Table 4, and those for the three-state model but with observation times no longer informative on the state are given in Table 5.
To calculate the marginal likelihood, integrating out the hidden Markov process, let $\mathbf{L}(t;\boldsymbol{\beta})$ be a diagonal matrix with $L_{k,k}=f(o_{t}|X_{\tau_{t}}=k)$. When the first event is guaranteed to occur at time 00, the full likelihood for a the set of event times and observations for a single person is:
$\mathcal{L}(\boldsymbol{\beta},\boldsymbol{\lambda},\mathbf{Q},\boldsymbol{\nu})\propto\boldsymbol{\nu}^{\top}\mathbf{L}(0;\boldsymbol{\beta})\exp[(\mathbf{Q}-\boldsymbol{\lambda})\tau_{1}]\left\{\prod_{t=1}^{T}\boldsymbol{\Lambda}\mathbf{L}(t;\boldsymbol{\beta})\exp[(\mathbf{Q}-\boldsymbol{\Lambda})(\tau_{t+1}-\tau_{t})]\right\}\mathbf{1},$
where $\tau_{T+1}=t_{e}$, the end of the observation window. For each of the three- and four-state CTMC-MMPP models we evaluate this expression at the model’s posterior mean parameter estimate. For the BIC calculations, the three-state model has 14 parameters and the four-state model has 23; the number of patients is $n=1000$.
Table 4: Posterior median of four-state MMPP parameters associated with 95% credible intervals.
Parameter State 1 State 2 State 3
$\lambda$ (Marked point process rate) 4.82 (4.70, 4.96) 5.51 (5.40, 5.62) 7.71 (7.54, 7.89)
$\nu$ (initial state distribution) 0.33 (0.30, 0.37) 0.35 (0.30, 0.40) 0.32 (0.28, 0.36)
Expected number of drugs dispensed (GP) 0.70 (0.66, 0.75) 4.25 (4.21, 4.28) 8.66 (8.62, 8.70)
Expected number of drugs dispensed (ED) 0.43 (0.37, 0.50) 4.01 (3.94, 4.08) 8.65 (8.58, 8.73)
Expected number of drugs dispensed (HOSP) 0.51 (0.45, 0.58) 3.81 (3.73, 3.88) 8.48 (8.39, 8.56)
Expected number of drugs dispensed (SPEC) 0.51 (0.45, 0.58) 4.15 (4.07, 4.22) 8.65 (8.56, 8.74)
$q_{1.}$ (Transition rate from State 1) -- 0.28 (0.24, 0.32) 0.02 (0.01, 0.04)
$q_{2.}$ (Transition rate from State 2) -- -- 0.18 (0.17, 0.20)
$q_{.4}$ (Transition rate to death state) 0.07 (0.06, 0.09) 0.07 (0.06, 0.08) 0.23 (0.21, 0.26)
Table 5: Posterior median of three-state CTHMM (with observation timing not informative on the state) parameters associated with 95% credible intervals.
Parameter State 1 State 2
$\nu$ (initial state distribution) 0.76 (0.73, 0.79) 0.24 (0.21, 0.27)
Expected number of drugs dispensed (GP) 3.59 (3.55, 3.62) 8.37 (8.33, 8.41)
Expected number of drugs dispensed (ED) 3.73 (3.65, 3.80) 8.81 (8.73, 8.90)
Expected number of drugs dispensed (HOSP) 3.32 (3.24, 3.40) 8.52 (8.43, 8.61)
Expected number of drugs dispensed (SPEC) 3.51 (3.43, 3.59) 8.45 (8.35, 8.55)
$q_{1.}$ (Transition rate from State 1) -- 0.16 (0.14, 0.18)
$q_{.3}$ (Transition rate to death state) 0.09 (0.08, 0.11) 0.04 (0.03, 0.05)